OpenStudy (aivantettet26):

find the first derivative of y = 4cos(sin3x)

OpenStudy (anonymous):

y=4cos(sin3x) Find the derivative of the expression. (d)/(dx) 4cos(sin3x) The chain rule states that the derivative of a composite function (f o g)' is equal to (f' o g)*g'. To find the derivative of 4cos(sin3x), find the derivatives of each portion of the function and use the chain rule formula. (d)/(du) 4cos(u)*(d)/(dx) sin3x The derivative of 4cos(u) is -(4sin(u)). (d)/(du) 4cos(u)=-(4sin(u)) Multiply -1 by the 4sin(u) inside the parentheses. (d)/(du) 4cos(u)=-4sin(u) The chain rule states that the derivative of a composite function (h o j)' is equal to (h' o j)*j'. To find the derivative of sin3x, find the derivatives of each portion of the function and use the chain rule formula. (d)/(du) sin(u)*(d)/(dx) 3x The derivative of sin(u) is cos(u). (d)/(du) sin(u)=cos(u) To find the derivative of 3x, multiply the base (x) by the exponent (1), then subtract 1 from the exponent (1-1=0). Since the exponent is now 0, x is eliminated from the term. (d)/(dx) 3x=3 Replace the variable u with 3x in the expression. (d)/(du) sin(u)=cos((3x)) Replace the variable u with 3x in the expression. 3 Form the derivative by substituting the values for each portion into the chain rule formula. =cos((3x))*3 Multiply cos((3x)) by 3 to get 3cos((3x)). (d)/(dx) sin3x=3cos((3x)) Replace the variable u with sin3x in the expression. (d)/(du) 4cos(u)=-4sin((sin3x)) Replace the variable u with sin3x in the expression. 3cos((3x)) Form the derivative by substituting the values for each portion into the chain rule formula. =-4sin((sin3x))*3cos((3x)) Multiply -4sin((sin3x)) by 3cos((3x)) to get -12sin((sin3x))cos((3x)). (d)/(dx) 4cos(sin3x)=-12sin((sin3x))cos((3x))

OpenStudy (aivantettet26):

that's a very long reply O.o. thank you for your efforts! my answer is -12cos(sin3x)sin(sin3x)cos(3x)

OpenStudy (aivantettet26):

but im not really sure

OpenStudy (bahrom7893):

\[y = 4Cos(Sin(3x))\]\[y'=4(-Sin(Sin(3x)))*(Sin(3x))'<-\space Chain \space Rule\]

OpenStudy (bahrom7893):

now just differentiate the second bit: \[(Sin(3x))' = Cos(3x)*(3x)'=Cos(3x)*3=3Cos(3x)\]

OpenStudy (bahrom7893):

Again I used the chain rule.

OpenStudy (bahrom7893):

So now putting the two equations together, we get: \[y'=4(-Sin(Sin(3x)))*(3Cos(3x))=-12Sin(Sin(3x))*Cos(3x)\]

OpenStudy (aivantettet26):

thank you! I saw my mistake my answer is -12sin(sin3x)cos(3x) and then i saw your reponse thank you for valiant effort!

OpenStudy (anonymous):

Yep looks good! Nice work @bahrom7893

OpenStudy (bahrom7893):

Cheers